# Temporal-Aware Evaluation and Learning for Temporal Graph Neural Networks

### v1

@@ 3 Preliminary and Background
The proof of Theorem 3.1 can be found in supplementary material. Theorem 3.1 demonstrates that instance-based evaluation metrics cannot differentiate predictions if the number of disagreements with the ground truth is the same. Essentially, such metrics reduce all diverse information (e.g., temporal information) of predictions to a mere disagreement count. This severely limits the expressiveness of these metrics, making them inadequate for capturing insightful information about predictions within the temporal process.
@@ 4 Methodology
Building on the previous discussion regarding the limitations of existing evaluation metrics, this section introduces a novel temporal-aware evaluation metric derived from the concept of Hopkins statistics (Banerjee and Dave 2004). Specifically, we focus on detecting volatility clusters within predictions, which have significant implications for algorithms and systems, as discussed earlier. Additionally, based on this proposed evaluation metric, we introduce a novel temporal-aware learning objective for TGNNs. The pseudo-code for the complete procedure is provided in the supplementary material.
@@
Given a test period $T_{\mathrm{test}}$ with an event sequence $\mathcal{E}_{\mathrm{test}}$, let $\mathbf{Y}$ and $\widehat{\mathbf{Y}}$ represent the ground truth and the predictions of the model on the test set, respectively. Let $\mathcal{E}_{\mathrm{disg}}$ denote the set of disagreement events with cardinality $k$. We first compute the sum of distances from the disagreement set to the test set as:
$\mathrm{D}_{\mathrm{disg}}$ $=\sum_{e\in\mathcal{E}_{\mathrm{disg}}}\mathrm{d}(e,\mathcal{E}_{\mathrm{test}}),$ (4.1)
$\mathrm{d}(e,\mathcal{E}_{\mathrm{test}})$ $=\min\left\{|t_{e}-t_{e^{\prime}}|\bigg{|}e^{\prime}\in\mathcal{E}_{\mathrm{test}},e^{\prime}\neq e\right\}.$ (4.2)
$\mathrm{d}(e,\mathcal{E}_{\mathrm{test}})$ calculates the time difference between event $e$ and the closest event in the test set. Then, $\mathrm{D}_{\mathrm{disg}}$ is a sum of such distances for the disagreement set. Next, we generate a set $\mathcal{E}_{\mathrm{r}}$ of $k$ events by randomly sampling from the test set $\mathcal{E}_{\mathrm{test}}$. Similarly, we compute its distance to the test set as:
$\mathrm{D}_{\mathrm{r}}$ $=\sum_{e\in\mathcal{E}_{\mathrm{r}}}\mathrm{d}(e,\mathcal{E}_{\mathrm{test}}).$ (4.3)
$\mathrm{D}_{\mathrm{r}}$ serves as a reference for the distance to the test set if the samples are randomly drawn. Then, we can compute the distance between the set $\mathcal{E}_{\mathrm{disg}}$ with respect to $\mathcal{E}_{\mathrm{test}}$ as:
$\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}})=\frac{\mathrm{D}_{\mathrm{r}}}{\mathrm{D}_{\mathrm{r}}+\mathrm{D}_{\mathrm{disg}}},$
where $\mathrm{D}_{\mathrm{r}}$ and $\mathrm{D}_{\mathrm{disg}}$ are described above. The formulation shows that $\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}})$ compares the temporal distance between predictions relative to random sampling. The ratio format confines the value within the range of 00 to $1$. The $\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}})$ statistic provides insights into the distribution of data points. If $\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}})$ is close to 00, it indicates that the data points are clustered, with the sum of distances from randomly generated points to their nearest neighbors being significantly smaller than that from the sampled data points. Conversely, if $\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}})$ is close to $1$, it suggests that the data points are well-spaced (dispersed), resulting in larger distances for randomly generated points compared to those from sampled data points. When $\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}})$ approximates $0.5$, it indicates a random distribution with no significant clustering or regular pattern, as both randomly generated points and sampled data points exhibit similar nearest neighbour distances.
@@
$\mathrm{VCS}=|1/2-\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}},\tau)|,$
$=\left|\frac{1}{2}-\frac{1}{\tau}\sum_{i=1}^{\tau}\frac{\mathrm{D}_{\mathrm{r}}^{(i)}}{\mathrm{D}_{\mathrm{r}}^{(i)}+\mathrm{D}_{\mathrm{disg}}}\right|.$
@@
$\mathcal{T}_{\mathrm{soft}}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}})=\frac{\widehat{\mathrm{D}}_{\mathrm{r}}}{\widehat{\mathrm{D}}_{\mathrm{r}}+\widehat{\mathrm{D}}_{\mathrm{disg}}},$
@@
^$\widehat{\mathcal{L}}(\widehat{\mathbf{Y}},\mathbf{Y})=\mathcal{L}(\widehat{\mathbf{Y}},\mathbf{Y})+\gamma\left\|\frac{1}{2}-\mathcal{T}_{\mathrm{soft}}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}})\right\|^{2},$ (4.4)
@@ 5 Empirical Study
We use five public dynamic graph benchmark datasets: Reddit, Wikipedia, MOOC, LastFM, and GDELT (Poursafaei et al. 2022). We evaluate six state-of-the-art TGNN models, with two models from each of the three categories of TGNNs mentioned: TGN (Rossi et al. 2021) & Tiger (Zhang et al. 2023) (memory-based TGNNs), TCL (Wang et al. 2021b) & TGAT (Xu et al. 2020a) (attention-based TGNNs), and JOIDE (Kumar, Zhang, and Leskovec 2019) & DyRep (Trivedi et al. 2019) (RNN-based TGNNs). We adopt the implementation of these baselines from (Zhou et al. 2022; Poursafaei et al. 2022; Huang et al. 2024). Detailed descriptions of the datasets and models are provided in the supplementary material for completeness.
@@ Appendix B Algorithm and Further Discussion
  Input: $\mathcal{E}_{\mathrm{test}}$, $\mathbf{Y}$, $\widehat{\mathbf{Y}}$
@@
$\mathrm{D}_{\mathrm{disg}}$ $=\sum_{e\in\mathcal{E}_{\mathrm{disg}}}\mathrm{d}(e,\mathcal{E}_{\mathrm{test}}),$
$\mathrm{d}(e,\mathcal{E}_{\mathrm{test}})$ $=\min\left\{|t_{e}-t_{e^{\prime}}|\bigg{|}e^{\prime}\in\mathcal{E}_{\mathrm{test}},e^{\prime}\neq e\right\}.$
  for i=1 to $\tau$ do
     sample a random set form $\mathcal{E}_{\mathrm{r}}^{(t)}$ of the same cardinality as $\mathcal{E}_{\mathrm{disg}}$.
@@
$\mathrm{VCS}=|1/2-\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}},\tau)|,$
$=\left|\frac{1}{2}-\frac{1}{\tau}\sum_{i=1}^{\tau}\frac{\mathrm{D}_{\mathrm{r}}^{(i)}}{\mathrm{D}_{\mathrm{r}}^{(i)}+\mathrm{D}_{\mathrm{disg}}}\right|.$
@@
$\mathcal{L}(\bar{B}_{i-1})+\gamma\cdot\left\|\frac{1}{2}-\mathcal{T}_{\mathrm{soft}}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}},\tau)\right\|.$

### v2

@@ 3 Preliminary and Background
footnotetext: Extra technical details such as proof, pseudo-code and experimental setting can be found in extended arxiv version (Su and Wu 2024).
Theorem 3.1 demonstrates that instance-based evaluation metrics cannot differentiate predictions if the number of disagreements with the ground truth is the same. Essentially, such metrics reduce all diverse information (e.g., temporal information) of predictions to a mere disagreement count. This severely limits the expressiveness of these metrics, making them inadequate for capturing insightful information about predictions within the temporal process.
@@ 4 Methodology
Building on the previous discussion regarding the limitations of existing evaluation metrics, this section introduces a novel temporal-aware evaluation metric derived from the concept of Hopkins statistics (Banerjee and Dave 2004). Specifically, we focus on detecting volatility clusters within predictions, which have significant implications for algorithms and systems, as discussed earlier. Additionally, based on this proposed evaluation metric, we introduce a novel temporal-aware learning objective for TGNNs.
@@
Given a test period $T_{\mathrm{test}}$, let $\mathbf{Y}$ and $\widehat{\mathbf{Y}}$ represent the ground truth and the predictions of the model on the test set, respectively. Let $\mathcal{E}_{\mathrm{disg}}$ denote the set of disagreement events with cardinality $K$ and let ^$\widehat{\mathcal{E}}_{\mathrm{disg}}$ denote $k<K$ samples from $\mathcal{E}_{\mathrm{disg}}$ We first compute the sum of distances from the sampled disagreement set to the disagreement as:
$\mathrm{D}_{\mathrm{disg}}$ $=\sum_{e\in\widehat{\mathcal{E}}_{\mathrm{disg}}}\mathrm{d}(e,\mathcal{E}_{\mathrm{disg}}),$ (4.1)
$\mathrm{d}(e,\mathcal{E}_{\mathrm{disg}})$ $=\min\left\{|t_{e}-t_{e^{\prime}}|\bigg{|}e^{\prime}\in\mathcal{E}_{\mathrm{disg}},e^{\prime}\neq e\right\}.$ (4.2)
$\mathrm{d}(e,\mathcal{E}_{\mathrm{disg}})$ calculates the time difference between event $e$ and the closest event in the given set. Then, $\mathrm{D}_{\mathrm{disg}}$ is a sum of such distances for the disagreement set. Next, we generate a set $\mathcal{E}_{\mathrm{r}}$ of $k$ events by uniformly randomly sampling from the test period $T_{\mathrm{test}}$. Similarly, we compute its distance to the disagreement as:
$\mathrm{D}_{\mathrm{r}}$ $=\sum_{e\in\mathcal{E}_{\mathrm{r}}}\mathrm{d}(e,\mathcal{E}_{\mathrm{disg}}).$ (4.3)
$\mathrm{D}_{\mathrm{r}}$ serves as a reference for the distance to the disagreement if the samples are randomly drawn. Then, we can compute relative statistics between the set $\mathcal{E}_{\mathrm{disg}}$ and $\mathcal{E}_{\mathrm{r}}$ as:
$\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{r}})=\frac{\mathrm{D}_{\mathrm{r}}}{\mathrm{D}_{\mathrm{r}}+\mathrm{D}_{\mathrm{disg}}},$
where $\mathrm{D}_{\mathrm{r}}$ and $\mathrm{D}_{\mathrm{disg}}$ are described above. The formulation shows that $\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{r}})$ compares the temporal distance between predictions relative to random sampling. The ratio format confines the value within the range of 00 to $1$. The $\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{r}})$ statistic provides insights into the distribution of data points. If $\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}})$ is close to $1$, it indicates that the data points are clustered, with the sum of distances from randomly generated points to their nearest neighbors being significantly larger than that from the sampled data points. Conversely, if $\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}})$ is close to 00, it could suggest that the data points are regularly-spaced, resulting in smaller distances for randomly generated points compared to those from sampled data points. When $\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{test}})$ approximates $0.5$, it indicates a random distribution with no significant clustering or regular pattern, as both randomly generated points and sampled data points exhibit similar nearest neighbour distances.
@@
$\mathrm{VCS}=|1/2-\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{r}},\tau)|,$
$=\left|\frac{1}{2}-\frac{1}{\tau}\sum_{i=1}^{\tau}\frac{\mathrm{D}_{\mathrm{r}}^{(i)}}{\mathrm{D}_{\mathrm{r}}^{(i)}+\mathrm{D}_{\mathrm{disg}}^{(i)}}\right|.$
@@
$\mathcal{T}_{\mathrm{soft}}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{r}})=\frac{\widehat{\mathrm{D}}_{\mathrm{r}}}{\widehat{\mathrm{D}}_{\mathrm{r}}+\widehat{\mathrm{D}}_{\mathrm{disg}}},$
@@
^$\widehat{\mathcal{L}}(\widehat{\mathbf{Y}},\mathbf{Y})=\mathcal{L}(\widehat{\mathbf{Y}},\mathbf{Y})+\gamma\left\|\frac{1}{2}-\mathcal{T}_{\mathrm{soft}}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{r}})\right\|^{2},$ (4.4)
@@ 5 Empirical Study
We use five public dynamic graph benchmark datasets: Reddit, Wikipedia, MOOC, LastFM, and GDELT (Poursafaei et al. 2022). We evaluate six state-of-the-art TGNN models, with two models from each of the three categories of TGNNs mentioned: TGN (Rossi et al. 2021) & Tiger (Zhang et al. 2023) (memory-based TGNNs), TCL (Wang et al. 2021b) & TGAT (Xu et al. 2020a) (attention-based TGNNs), and JOIDE (Kumar, Zhang, and Leskovec 2019) & DyRep (Trivedi et al. 2019) (RNN-based TGNNs). We adopt the implementation of these baselines from (Zhou et al. 2022; Poursafaei et al. 2022; Huang et al. 2024).
@@ Appendix B Algorithm and Further Discussion
  Input: $\mathbf{Y}$, $\widehat{\mathbf{Y}}$
@@
$\mathrm{D}_{\mathrm{disg}}^{(i)}$ $=\sum_{e\in\widehat{\mathcal{E}}_{\mathrm{disg}}^{(i)}}\mathrm{d}(e,\mathcal{E}_{\mathrm{disg}}),$
$\mathrm{d}(e,\mathcal{E}_{\mathrm{disg}})$ $=\min\left\{|t_{e}-t_{e^{\prime}}|\bigg{|}e^{\prime}\in\mathcal{E}_{\mathrm{disg}},e^{\prime}\neq e\right\}.$
     generate a random set form $\mathcal{E}_{\mathrm{r}}^{(i)}$ of the same cardinality as ^$\widehat{\mathcal{E}}_{\mathrm{disg}}^{(i)}$.
@@
$\mathrm{VCS}=|1/2-\mathcal{T}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{r}},\tau)|,$
$=\left|\frac{1}{2}-\frac{1}{\tau}\sum_{i=1}^{\tau}\frac{\mathrm{D}_{\mathrm{r}}^{(i)}}{\mathrm{D}_{\mathrm{r}}^{(i)}+\mathrm{D}_{\mathrm{disg}}^{(i)}}\right|.$
@@
$\mathcal{L}(\bar{B}_{i-1})+\gamma\cdot\left\|\frac{1}{2}-\mathcal{T}_{\mathrm{soft}}(\mathcal{E}_{\mathrm{disg}},\mathcal{E}_{\mathrm{r}},\tau)\right\|.$